# Comparison of 4.5PN and 2SF gravitational energy fluxes from quasicircular compact binaries

### v1

@@ IV Comparison
In Fig. 1 we plot the GSF data at 2SF order, i.e. corresponding to the $\mathcal{O}(\nu)$ terms in the normalized flux (43), together with the PN predictions at various orders up to 4.5PN. We nicely find that both the 4PN and 4.5PN are very close to the 2SF prediction, say until the relativistic regime at around $x\sim 0.15$. Previous PN approximations do clearly less well. While 3PN is still rather good, the half-integer (odd-parity) 3.5PN is off. By contrast, 2.5PN is surprisingly close to the 2SF result, while the even-parity 2PN is the worst approximation in the series. We also note that 1PN does relatively well. Here the PN expansion is in Taylor-expanded form, without resummation techniques applied.
Figure 2 goes further in the comparison by checking that the PN behaviour of the flux is indeed contained in and reproduced by the 2SF result for various orders. Observe in particular the residual obtained after subtraction of all known PN terms from the 2SF result. The residual is consistent with the expected 5PN behaviour $\mathcal{O}(x^{5})$ of the systematic PN error. [^4] There should be also some logarithmic terms $\mathcal{O}(x^{5}\ln x)$, but one cannot distinguish between the slopes of $x^{5}$ and $x^{5}\ln x$ on the scale of the plot. For small values of $x$ the non-smoothness of the residual after subtracting the 4PN and 4.5PN results stems from numerical error in the GSF calculation. The level of this noise is consistent with the error in each $(\ell,m)$ being similar to the estimated error in the (2,2) mode as shown in Fig. 7 of [5].
Although Fig. 2 provides convincing evidence that the 2SF and 4.5PN results agree for the total Newtonian-normalized flux $\hat{\mathcal{F}}$, we also find compelling evidence that the individual mode contributions ^$\hat{\mathcal{{F}}}_{\ell m}$ do not agree beyond 3.5PN order. We explore that comparison in Appendix A. This disagreement at the level of individual mode contributions, which cancels when the contributions are summed, might indicate that the GSF and PN waveforms are computed in different asymptotic Bondi-Metzner-Sachs (BMS) frames [48].
@@
In Fig. 3 we assess the accuracy and convergence of the SF fluxes when benchmarked against PN and numerical relativity (NR); this complements the analogous assessment of PN in Fig. 1. Here we include the normalized total flux at 1SF [i.e., neglecting $\mathcal{O}(\nu)$] and 2SF [neglecting $\mathcal{O}(\nu^{2})$] alongside 4.5PN and NR results for two mass ratios: $q=1$ (left panel) and $q=10$ (right panel). We see that for both mass ratios, the 2SF flux provides a marked improvement over the 1SF result. The SF expansion, moving from 1SF to 2SF, converges well toward the NR result, achieving nice agreement with NR for the large mass ratio $q=10$ throughout the strong field regime $x\gtrsim 0.07$ where the NR data is available. The SF expansion also converges nicely toward the PN result in the weak field, with the PN and 2SF results agreeing extremely well in the mildly relativistic regime, say $x\lesssim 0.07$, for both mass ratios. To further illustrate the convergence, we also include curves for the 4.5PN flux truncated at 1SF and at 2SF. The latter agrees extremely well with the full 4.5PN flux for all values of $x$, showing that higher-order corrections $\mathcal{O}(\nu^{2})$ are numerically small even at $q=1$. We note that for $q=1$, the 4.5PN flux remains more accurate than the 2SF flux, when measured against NR, even in the strong field up to $x\sim 0.1$ (a very relativistic value by PN standards). Figure 3 thus illustrates the importance of synergies between the PN, SF, and NR approaches to cover the whole range of scenarios of compact binary coalescences.
@@ V Conclusion
Previous comparisons of the 4.5PN quasicircular flux were performed using the GSF flux to first order in the mass ratio. To 1SF level the GSF results can be derived analytically to high order in the PN expansion [21, 22, 49, 50], and the PN coefficients are perfectly consistent up to 4.5PN order with direct PN calculations valid for arbitrary mass ratios. Note that at 1SF/0PA order, the flux at infinity for quasicircular inspirals is identical to the flux generated by a particle on a fixed background geodesic.
The new comparisons at the 2SF level reported in this paper are numerical, building on the earlier comparisons with the 3.5PN flux in Ref. [5]. The 2SF fluxes take into account the contribution coming from the 1SF deviation in the motion of the particle generating the GW and flux at infinity, as well as quadratic nonlinearities in waveform generation and propagation. As shown in Figs. 1, 2, and 3, the agreement with the PN coefficients for the total flux is excellent.
A key aspect of our analysis is that invariant comparisons require precise knowledge of how the orbital frequency relates to the waveform frequency. Agreement between the 4.5PN and 2SF fluxes is only achieved when the fluxes are computed as functions of an invariant waveform frequency, eliminating the gauge/slicing dependence of the relationship between the orbit and the waveform.
Another interesting outcome of our comparison is that while the total PN and SF fluxes agree through 4.5PN order, they disagree at the level of individual spherical-harmonic mode contributions beyond 3.5PN order. This is highly suggestive of the two waveforms being in different BMS frames. Future work will explore the choice of frame in the 2SF waveform and determine whether mode-by-mode agreement with PN can be achieved through a frame transformation, as has been done in numerical relativity [51].
@@ VI Acknowledgments
N.W. acknowledges support from a Royal Society – Science Foundation Ireland University Research Fellowship. This publication has emanated from research conducted with the financial support of Science Foundation Ireland under grant number 22/RS-URF-R/3825. D.T. received support from the Czech Academy of Sciences under the grant number LQ100102101. A.P. acknowledges the support of a Royal Society University Research Fellowship and the UKRI Frontier Research Grant GWModels, as selected by the ERC and funded by UKRI under the Horizon Europe Guarantee scheme [grant number EP/Y008251/1].
Appendix A Comparison of individual mode contributions to the flux
The comparisons in the body of the paper are for the total flux to infinity, summed over $(\ell,m)$ mode contributions. However, additional information about the PN and SF waveform amplitudes can be gleaned by comparing the individual $(\ell,m)$ contributions.
Figure 4: (Left panel) Comparison between $\mathcal{O}(\nu)$ GSF and PN results for the (Newtonian-normalized) fluxes for the (2,2) mode. At the level of individual modes we do not see agreement after 3.5PN. (Right panel) Same as the left panel but for the (3,3) mode. The PN results for the (2,2) and (3,3) modes are respectively given by (47a) and (47c).
At 4PN order, the modes can be straightforwardly obtained by inserting the $(2,2)$ mode amplitude (II) and Eqs. (3.4) of [36] for the other modes, into Eq. (7). Moreover, we can extend this result at 4.5PN order with the sole knowledge of the hereditary part of the first time derivative of the radiative moments [14]. Using (2.6)–(2.7) of [29] it comes
@@ Appendix A Comparison of individual mode contributions to the flux
The explicit 4.5PN expressions of the decomposition of the total flux are given in Eq. (47) hereafter; we have verified that these expressions agree perfectly with Eqs. (A.3-27) of [22] at first order in the mass ratio. We present the results in terms of the normalized fluxes ^$\hat{\mathcal{F}}_{\ell m}$ defined in Eq. (43).
Finally we present in Figs. 4 the comparison between the 2SF and PN results for the flux contributions from the individual spherical-harmonic modes (2,2) and (3,3). We find that the individual modes do not agree well beyond 3.5PN order, even though the sum of the modes agrees as we have seen in Fig. 2. This disagreement on the separate modes may be due to the fact that the GSF waveform has been computed in an asymptotic frame different from that of the PN waveform.
A difference in asymptotic frame would not be surprising. Historically, little work has been done to control the BMS frame in GSF calculations, which have mostly used intuition from the Newtonian-order analysis in Ref. [52]. A BMS transformation, whether an ordinary Poincaré transformation or a supertranslation, can significantly alter the individual $(\ell,m)$ modes of the waveform [48, 51], and this freedom remains unexplored in GSF calculations. An analysis of this freedom, and whether it can resolve the mode-by-mode discrepancy we find, is left for future work.
^$\hat{\mathcal{F}}_{22}$ $=1+\left(-\frac{107}{21}+\frac{55}{21}\nu\right)x+4\pi x^{3/2}+\left(\frac{4784}{1323}-\frac{87691}{5292}\nu+\frac{5851}{1323}\nu^{2}\right)x^{2}+\left(-\frac{428}{21}+\frac{178}{21}\nu\right)\pi x^{5/2}$
@@
^$\hat{\mathcal{F}}_{21}$ $=\left(\frac{1}{36}-\frac{\nu}{9}\right)x+\left(-\frac{17}{504}+\frac{11}{63}\nu-\frac{10}{63}\nu^{2}\right)x^{2}+\left(\frac{1}{18}-\frac{2}{9}\nu\right)\pi x^{5/2}$
@@
^$\hat{\mathcal{F}}_{33}$ $=\left(\frac{1215}{896}-\frac{1215}{224}\nu\right)x+\left(-\frac{1215}{112}+\frac{10935}{224}\nu-\frac{1215}{56}\nu^{2}\right)x^{2}+\left(\frac{3645}{448}-\frac{3645}{112}\nu\right)\pi x^{5/2}$
@@
^$\hat{\mathcal{F}}_{32}$ $=\left(\frac{5}{63}-\frac{10}{21}\nu+\frac{5}{7}\nu^{2}\right)x^{2}+\left(-\frac{193}{567}+\frac{1304}{567}\nu-\frac{2540}{567}\nu^{2}+\frac{365}{189}\nu^{3}\right)x^{3}+\left(\frac{20}{63}-\frac{40}{21}\nu+\frac{20}{7}\nu^{2}\right)\pi x^{7/2}$
@@
^$\hat{\mathcal{F}}_{31}$ $=\left(\frac{1}{8064}-\frac{\nu}{2016}\right)x+\left(-\frac{1}{1512}+\frac{5}{2016}\nu+\frac{\nu^{2}}{1512}\right)x^{2}+\left(\frac{1}{4032}-\frac{\nu}{1008}\right)\pi x^{5/2}$
@@
^$\hat{\mathcal{F}}_{44}$ $=\left(\frac{1280}{567}-\frac{2560}{189}\nu+\frac{1280}{63}\nu^{2}\right)x^{2}+\left(-\frac{151808}{6237}+\frac{2995712}{18711}\nu-\frac{1853440}{6237}\nu^{2}+\frac{32000}{297}\nu^{3}\right)x^{3}$
@@
^$\hat{\mathcal{F}}_{43}$ $=\left(\frac{729}{4480}-\frac{729}{560}\nu+\frac{729}{224}\nu^{2}-\frac{729}{280}\nu^{3}\right)x^{3}+\Bigg{(}-\frac{28431}{24640}+\frac{198045}{19712}\nu-\frac{288441}{9856}\nu^{2}+\frac{20169}{616}\nu^{3}-\frac{31833}{3080}\nu^{4}\Bigg{)}x^{4}$
@@
^$\hat{\mathcal{F}}_{42}$ $=\left(\frac{5}{3969}-\frac{10}{1323}\nu+\frac{5}{441}\nu^{2}\right)x^{2}+\left(-\frac{437}{43659}+\frac{7958}{130977}\nu-\frac{4120}{43659}\nu^{2}+\frac{95}{14553}\nu^{3}\right)x^{3}$
@@
^$\hat{\mathcal{F}}_{41}$ $=\left(\frac{1}{282240}-\frac{\nu}{35280}+\frac{\nu^{2}}{14112}-\frac{\nu^{3}}{17640}\right)x^{3}+\Bigg{(}-\frac{101}{4656960}+\frac{229}{1241856}\nu-\frac{107}{206976}\nu^{2}+\frac{\nu^{3}}{1848}-\frac{83}{582120}\nu^{4}\Bigg{)}x^{4}$
@@
^$\hat{\mathcal{F}}_{55}$ $=\left(\frac{9765625}{2433024}-\frac{9765625}{304128}\nu+\frac{48828125}{608256}\nu^{2}-\frac{9765625}{152064}\nu^{3}\right)x^{3}$
@@
^$\hat{\mathcal{F}}_{54}$ $=\Bigg{(}\frac{4096}{13365}-\frac{8192}{2673}\nu+\frac{28672}{2673}\nu^{2}-\frac{40960}{2673}\nu^{3}+\frac{20480}{2673}\nu^{4}\Bigg{)}x^{4}+\mathcal{O}(x^{5})$ (47k)
^$\hat{\mathcal{F}}_{53}$ $=\left(\frac{2187}{450560}-\frac{2187}{56320}\nu+\frac{2187}{22528}\nu^{2}-\frac{2187}{28160}\nu^{3}\right)x^{3}$
@@
^$\hat{\mathcal{F}}_{52}$ $=\Bigg{(}\frac{4}{40095}-\frac{8}{8019}\nu+\frac{28}{8019}\nu^{2}-\frac{40}{8019}\nu^{3}+\frac{20}{8019}\nu^{4}\Bigg{)}x^{4}+\mathcal{O}(x^{5})$ (47m)
^$\hat{\mathcal{F}}_{51}$ $=\left(\frac{1}{127733760}-\frac{\nu}{15966720}+\frac{\nu^{2}}{6386688}-\frac{\nu^{3}}{7983360}\right)x^{3}$

### v2

@@ IV Comparison
In Fig. 1 we plot the GSF data for $\hat{\mathcal{F}}^{(2)}=\mathcal{F}^{(2)}/\mathcal{F}_{\mathrm{N}}$, i.e. corresponding to the terms in the normalized flux (43) that are subleading in the mass ratio, together with the PN predictions at various orders, up to 4.5PN. We find that both the 4PN and 4.5PN are numerically very close to the 2SF prediction until the relativistic regime at around $x\simeq 0.15$. For $x\lesssim 0.12$, the 4.5PN series is in better agreement with the numerical data than the 4PN series; but around $x\simeq 0.15$, it is the opposite. Lower-order PN approximations clearly agree less well. While the 3PN prediction is still rather good, the half-integer (odd-parity) 3.5PN is off. By contrast, 2.5PN is surprisingly close to the 2SF result, while the even-parity 2PN is the worst approximation in the series. We also note that 1PN does relatively well. Here the PN expansion is in Taylor-expanded form, without resummation techniques applied.
Figure 2 goes further in the comparison by checking the residuals between the GSF data and PN series. We find that the residual obtained by subtracting the 3.5PN result from the numerical GSF data clearly follows the 4PN term. When further subtracting the 4PN series from the numerical GSF data, we observe that the amplitude of the residual is subdominant (with respect to the 4PN term) and close to the 4.5PN term. However, over the range of frequencies where we have GSF data, we cannot ascertain the slope nor conclusively say that the residual is tending to the 4.5PN result. Finally, after subtracting the 4.5PN term from the numerical GSF data, we again observe that the amplitude of the residual is subdominant but again, we cannot ascertain how the residual scales with $x$. Our inability to conclusively determine the scaling of the 4PN and 4.5PN residual stems from the limited range of frequencies over which we have accurate numerical GSF data. As the PN series is an asymptotic expansion, we only expect agreement with the numerical data for sufficiently small values of $x$. How small $x$ has to be depends on the PN residual being considered. In particular, the scaling of the residual will only become apparent for values of $x$ below which there are no more zero crossings. For example, in Fig. 2, there is a zero crossing in the 4.5PN residual at $x\simeq 0.11$. These zero crossings can in fact occur at much smaller values of $x$ than we have access to; we give an illustration of this by studying equivalent plots for the first-order flux in Appendix C. Thus, firmly establishing the scaling of the 4PN and 4.5PN residuals would require accurate GSF data at much smaller values of $x$ than we currently have access to. The level of this noise is consistent with the error in each $(\ell,m)$ being similar to the estimated error in the (2,2) mode as shown in Fig. 7 of [5].
Although the comparison with the total flux in Fig. 2 provides convincing evidence of agreement between the GSF and PN results, the precise scaling of the residual with 4PN and 4.5PN cannot be clearly established. Fortunately, we can build further confidence in our results by making comparisons at the level of individual $(\ell,m)$ modes, because each PN order and each $(\ell,m)$ mode exhibits different zero crossings. The comparison with the $(2,2)$ mode is shown in Fig. 3. For this mode, the comparison with the 3.5PN term pushes the residual into the numerical noise of the GSF data at small values of $x$. There is also evidence that there must be a further zero crossing in the 4PN residual at smaller values of $x$ than we have access to. The comparison with the $(3,3)$ and $(3,2)$ modes is much less noisy, even down to the 4.5PN residuals, as shown in Fig. 5. Similarly, good agreement is seen for many other modes – see the figures in Appendix B. We note here that agreement between the $(\ell,m)$ modes is not guaranteed unless the waveforms are computed in the same asymptotic Bondi-Metzner-Sachs (BMS) frames [48]. At present, in second-order GSF calculations, this is not well understood and requires more work; we briefly discuss this further in Appendix B.
Figure 5: (Left Panel) Comparison between the (Newtonian-normalized) GSF and PN total flux at different orders in the mass ratio expansion for $q=1$ ($\nu=0.25$). It is interesting to note that the 4.5PN expansion, including all orders in $\nu$ (black, dashed curve), is closely approximated by the PN expansion truncated at $\mathcal{O}(\nu^{2})$ (yellow, solid curve), despite the equal mass ratio. The two gray curves show the flux computed from an NR simulation by the SXS collaboration [49], specifically SXS:BBH:1132 [50]. The two gray curves are computed using, respectively, 3rd and 4th order polynomials to extrapolate the NR waveform from the edge of the computational domain to null infinity and thus the shaded gray region between them is an estimate on the numerical error in the simulation. (Right Panel) The same as the left panel but for $q=10$ ($\nu=0.0826$). The NR simulation in this plot is SXS:BBH:1107 [51].
@@
In Fig. 5 we assess the accuracy and convergence of the SF fluxes when benchmarked against PN and numerical relativity (NR); this complements the analogous assessment of PN in Fig. 1. Here we include the normalized total flux at 1SF [i.e., neglecting $\mathcal{O}(\nu)$] and 2SF [neglecting $\mathcal{O}(\nu^{2})$] alongside 4.5PN and NR results for two mass ratios: $q=1$ (left panel) and $q=10$ (right panel). We see that for both mass ratios, the 2SF flux provides a marked improvement over the 1SF result. The SF expansion, moving from 1SF to 2SF, converges well toward the NR result, achieving nice agreement with NR for the large mass ratio $q=10$ throughout the strong field regime $x\gtrsim 0.07$ where the NR data is available. The SF expansion also converges nicely toward the PN result in the weak field, with the PN and 2SF results agreeing extremely well in the mildly relativistic regime, say $x\lesssim 0.07$, for both mass ratios. To further illustrate the convergence, we also include curves for the 4.5PN flux truncated at 1SF and at 2SF. The latter agrees extremely well with the full 4.5PN flux for all values of $x$, showing that higher-order corrections $\mathcal{O}(\nu^{2})$ are numerically small even at $q=1$. We note that for $q=1$, the 4.5PN flux remains more accurate than the 2SF flux, when measured against NR, even in the strong field up to $x\simeq 0.1$ (a very relativistic value by PN standards). Figure 5 thus illustrates the importance of synergies between the PN, SF, and NR approaches to cover the whole range of scenarios of compact binary coalescences.
@@ V Conclusion
To 1SF level, the GSF flux has been derived analytically to high order in the PN expansion [21, 22, 52, 53], and the PN coefficients are perfectly consistent up to 4.5PN order with direct PN calculations valid for arbitrary mass ratios. Note that at 1SF/0PA order, the flux at infinity for quasicircular inspirals is identical to the flux generated by a particle on a fixed background geodesic.
The new comparisons at the 2SF level reported in this paper are numerical, building [^6] The second-order self-force data used in Ref. [5] contained an error that was numerically small but had a significant impact on some of the PN comparisons in that paper. An erratum giving full details is currently in preparation. The same error affected an earlier version of this paper [arXiv:2407.00366v1]. Using the incorrect data, we found good agreement at the level of the total flux, but stark disagreement at the level of individual modes. We interpreted this as mode-mixing, and suggested that it could arise from differences in the asymptotic frames used in the two calculations. Correcting this error in the data removed this apparent mode-mixing; we have therefore revised some of our conclusions. on earlier comparisons with the 3.5PN flux in Ref. [5]. The 2SF fluxes take into account the contribution coming from the 1SF deviation in the motion of the particle generating the GWs and flux at infinity, as well as quadratic nonlinearities in waveform generation and propagation.
Over the considered range of frequencies ($0.3<x<0.16$), we found that the 4PN and 4.5PN series agree much better with the numerical SF plot that any other lower-order PN truncation, see Fig. 1. We then studied the weak-field scaling of the residuals between the SF data and PN series at different orders, and found that the numerical data is compatible with the PN coefficients for the total flux, although we cannot conclusively demonstrate that the residual with the 4PN and 4.5PN results scale as expected over the range of frequencies for which we have GSF data. This is most likely due to the presence of zero crossings, which indicate that we have not yet reached the asymptotic regime where the comparison can be performed. Reaching this regime would require reducing the numerical noise in the 2SF data to levels significantly lower than what is currently achievable. To avoid any doubt, this does not mean that the numerical and PN fluxes disagree: one should keep in mind that near a zero crossing, the truncated PN series happen to agree numerically with the GSF data better that what is predicted by the asymptotic PN scaling. Finally, further comparisons at the level of the individual $(\ell,m)$ modes add further weight to the agreement, since the residuals for some $(\ell,m)$ modes do scale in the expected way – see Figs. 3, 5, 6, and 7.
Finally, note that a key aspect of our analysis is that invariant comparisons require precise knowledge of how the orbital frequency relates to the waveform frequency. Agreement between the 4.5PN and 2SF fluxes is only achieved when the fluxes are computed as functions of an invariant waveform frequency, eliminating the gauge/slicing dependence of the relationship between the orbit and the waveform [36].
@@ VI Acknowledgments
We are indebted to François Larrouturou for initiating this work with us, and for interesting discussions at an early stage of the project. A.P. is grateful to Thibault Damour for collaboration on a related project that uncovered an error in the 2SF flux data. N.W. acknowledges support from a Royal Society – Science Foundation Ireland University Research Fellowship. This publication has emanated from research conducted with the financial support of Science Foundation Ireland under grant number 22/RS-URF-R/3825. D.T. received support from the Czech Academy of Sciences under the grant number LQ100102101. A.P. acknowledges the support of a Royal Society University Research Fellowship and the UKRI Frontier Research Grant GWModels, as selected by the ERC and funded by UKRI under the Horizon Europe Guarantee scheme [grant number EP/Y008251/1]. This work makes use of the Black Hole Perturbation Toolkit.
Appendix A PN expansions for individual mode contributions to the flux
Some of the comparisons in this paper involve individual $(\ell,m)$ modes of the flux. As these were not given previously in the literature, we give them explicitly here. At 4PN order, the modes can be straightforwardly obtained by inserting the $(2,2)$ mode amplitude (II) and Eqs. (3.4) of [37] for the other modes, into Eq. (7). Moreover, we can extend this result at 4.5PN order with the sole knowledge of the hereditary part of the first time derivative of the radiative moments [14]. Using (2.6)–(2.7) of [29] it becomes
@@ Appendix A PN expansions for individual mode contributions to the flux
The explicit 4.5PN expressions of the decomposition of the total flux are given in Eq. (47) hereafter; we have verified that these expressions agree perfectly with Eqs. (A.3-27) of [22] at first order in the mass ratio. We present the results in terms of the normalized fluxes $\hat{\mathcal{F}}_{\ell m}$ defined in Eq. (43).
$\hat{\mathcal{F}}_{22}$ $=1+\left(-\frac{107}{21}+\frac{55}{21}\nu\right)x+4\pi x^{3/2}+\left(\frac{4784}{1323}-\frac{87691}{5292}\nu+\frac{5851}{1323}\nu^{2}\right)x^{2}+\left(-\frac{428}{21}+\frac{178}{21}\nu\right)\pi x^{5/2}$
@@
$\hat{\mathcal{F}}_{21}$ $=\left(\frac{1}{36}-\frac{\nu}{9}\right)x+\left(-\frac{17}{504}+\frac{11}{63}\nu-\frac{10}{63}\nu^{2}\right)x^{2}+\left(\frac{1}{18}-\frac{2}{9}\nu\right)\pi x^{5/2}$
@@
$\hat{\mathcal{F}}_{33}$ $=\left(\frac{1215}{896}-\frac{1215}{224}\nu\right)x+\left(-\frac{1215}{112}+\frac{10935}{224}\nu-\frac{1215}{56}\nu^{2}\right)x^{2}+\left(\frac{3645}{448}-\frac{3645}{112}\nu\right)\pi x^{5/2}$
@@
$\hat{\mathcal{F}}_{32}$ $=\left(\frac{5}{63}-\frac{10}{21}\nu+\frac{5}{7}\nu^{2}\right)x^{2}+\left(-\frac{193}{567}+\frac{1304}{567}\nu-\frac{2540}{567}\nu^{2}+\frac{365}{189}\nu^{3}\right)x^{3}+\left(\frac{20}{63}-\frac{40}{21}\nu+\frac{20}{7}\nu^{2}\right)\pi x^{7/2}$
@@
$\hat{\mathcal{F}}_{31}$ $=\left(\frac{1}{8064}-\frac{\nu}{2016}\right)x+\left(-\frac{1}{1512}+\frac{5}{2016}\nu+\frac{\nu^{2}}{1512}\right)x^{2}+\left(\frac{1}{4032}-\frac{\nu}{1008}\right)\pi x^{5/2}$
@@
$\hat{\mathcal{F}}_{44}$ $=\left(\frac{1280}{567}-\frac{2560}{189}\nu+\frac{1280}{63}\nu^{2}\right)x^{2}+\left(-\frac{151808}{6237}+\frac{2995712}{18711}\nu-\frac{1853440}{6237}\nu^{2}+\frac{32000}{297}\nu^{3}\right)x^{3}$
@@
$\hat{\mathcal{F}}_{43}$ $=\left(\frac{729}{4480}-\frac{729}{560}\nu+\frac{729}{224}\nu^{2}-\frac{729}{280}\nu^{3}\right)x^{3}+\Bigg(-\frac{28431}{24640}+\frac{198045}{19712}\nu-\frac{288441}{9856}\nu^{2}+\frac{20169}{616}\nu^{3}-\frac{31833}{3080}\nu^{4}\Bigg)x^{4}$
@@
$\hat{\mathcal{F}}_{42}$ $=\left(\frac{5}{3969}-\frac{10}{1323}\nu+\frac{5}{441}\nu^{2}\right)x^{2}+\left(-\frac{437}{43659}+\frac{7958}{130977}\nu-\frac{4120}{43659}\nu^{2}+\frac{95}{14553}\nu^{3}\right)x^{3}$
@@
$\hat{\mathcal{F}}_{41}$ $=\left(\frac{1}{282240}-\frac{\nu}{35280}+\frac{\nu^{2}}{14112}-\frac{\nu^{3}}{17640}\right)x^{3}+\Bigg(-\frac{101}{4656960}+\frac{229}{1241856}\nu-\frac{107}{206976}\nu^{2}+\frac{\nu^{3}}{1848}-\frac{83}{582120}\nu^{4}\Bigg)x^{4}$
@@
$\hat{\mathcal{F}}_{55}$ $=\left(\frac{9765625}{2433024}-\frac{9765625}{304128}\nu+\frac{48828125}{608256}\nu^{2}-\frac{9765625}{152064}\nu^{3}\right)x^{3}$
@@
$\hat{\mathcal{F}}_{54}$ $=\Bigg(\frac{4096}{13365}-\frac{8192}{2673}\nu+\frac{28672}{2673}\nu^{2}-\frac{40960}{2673}\nu^{3}+\frac{20480}{2673}\nu^{4}\Bigg)x^{4}+\mathcal{O}(x^{5})$ (47k)
$\hat{\mathcal{F}}_{53}$ $=\left(\frac{2187}{450560}-\frac{2187}{56320}\nu+\frac{2187}{22528}\nu^{2}-\frac{2187}{28160}\nu^{3}\right)x^{3}$
@@
$\hat{\mathcal{F}}_{52}$ $=\Bigg(\frac{4}{40095}-\frac{8}{8019}\nu+\frac{28}{8019}\nu^{2}-\frac{40}{8019}\nu^{3}+\frac{20}{8019}\nu^{4}\Bigg)x^{4}+\mathcal{O}(x^{5})$ (47m)
$\hat{\mathcal{F}}_{51}$ $=\left(\frac{1}{127733760}-\frac{\nu}{15966720}+\frac{\nu^{2}}{6386688}-\frac{\nu^{3}}{7983360}\right)x^{3}$